# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

### v2

@@ 5.2 Prediction Experiments
We compare COD against three representative prediction methods:
@@
    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws.
@@
    Passrate-compute prediction [1, 20]: Estimates large model passrates from smaller model passrates. We conduct 100 trials per evaluation set for smaller models to enhance reliability and report absolute prediction error on the passrate metric.
@@
Table 1: Absolute prediction error (in percentage points) on evaluation sets. Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red). $\downarrow$ indicates lower is better.
@@
    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within each cluster, and aggregates metrics, bypassing the subset-to-full mapping stage.
@@
Comparative results are shown in Table 1. Prediction accuracy is measured by the absolute error between predicted and actual performance. We report mean and max prediction errors across all evaluation sets, as well as errors for individual sets. Our complete COD approach significantly outperforms existing methods in both mean (1.68%) and maximum (2.38%) prediction errors, offering reliable guidance for large model training. While baselines perform well on some sets, their substantial errors on others undermine overall credibility.
Figure 4: Performance-compute relationship for different prediction methods on BBH, MATH, and MMLU-pro evaluation sets.
Fig. 4 visualizes the performance-compute relationships. On BBH, while all methods give similar final estimates, end-to-end and loss-intermediate methods fit small model data poorly. COD reveals a more complex, better-fitted multi-phase trajectory. On MATH and MMLU-pro, where predicting accelerated growth versus plateaus is crucial, the loss-intermediate method underestimates model ceilings, and the end-to-end method exceeds 3% error. COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets (e.g., MATH) and capture diminishing returns in saturated sets.
The COD method is also capable of predicting LLMs’ performance after continual training phase on high-quality data. We conduct experiments and analysis on continual-trained LLMs and report prediction results in Appendix D.
@@ 5.3.1 Comparison of Clustering Methods
We assess the impact of different clustering algorithms on prediction accuracy. The goal is to achieve tight intra-cluster difficulty similarity (low average distance to center) while maintaining cluster stability (min. 10 samples/cluster). We compare our Improved-MeanShift with DBScan, MeanShift, and K-Means. For K-Means, we adjust it to approximate our goals: (1) search for $k$ yielding min. cluster sizes $\approx 10$; (2) treat samples outside a radius (e.g., 2x average intra-cluster distance) from any cluster center as outliers, ensuring clusters don’t drop below 10 samples. We term this "Improved-KMeans" for this comparison. Clustering quality is measured by Intra-cluster Average Distance (IAD) and Outlier Rate (OR) (Table 2). Prediction benefits are measured by Extrapolation Error (EE) on the predictable subset and Final prediction Error (FE) (Table 3).
@@
K-Means 0.3236 - 0.2238 - 0.2238 - 0.6284 -
DBScan 0.4242 0.56 0.5131 0.53 0.4775 0.68 0.7113 18.92
MeanShift 0.2859 0.39 0.2852 0.61 0.2110 1.44 0.2679 20.72
Improved-KMeans 0.1609 2.85 0.1321 2.73 0.0902 2.22 0.1953 37.23
Improved-MeanShift 0.2225 4.40 0.1854 4.93 0.1463 2.66 0.2143 33.58
[…]
Method Mean Max MMLU-pro GSM8k MATH BBH
[…]
K-Means 3.62 3.76 8.16 8.99 3.69 3.69 0.01 0.00 2.62 2.34 8.16 8.99
DBScan 3.93 4.08 4.38 4.36 3.72 3.69 4.08 4.12 4.38 4.16 3.53 4.36
MeanShift 2.12 1.68 3.15 3.08 3.15 3.08 0.67 0.74 2.55 2.26 2.12 0.65
Improved-KMeans 1.33 1.84 3.92 4.08 0.56 0.61 3.92 4.08 0.81 0.51 0.02 2.17
Improved-MeanShift 1.23 1.66 2.20 2.23 1.27 1.35 2.20 2.23 1.14 1.28 0.29 1.77
Table 2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints. Table 3 confirms these methods lead to lower EE and FE. Although Improved-KMeans has the best IAD, its GSM8k prediction is poor. This is likely because K-Means requires pre-specifying $k$, and our search for $k$ can be unstable, leading to large errors on some sets. In contrast, our Improved-MeanShift, which automatically determines $k$ based on distance constraints, offers more stable clustering and the lowest maximum prediction error. Additional results in Sec. E.1 support these conclusions.
@@ 5.3.2 Different Extrapolation Formulas
Table 4: Ablation study of extrapolation formulas. EE, TR, FE shown for BBH, MATH, MMLU-pro.
@@
We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: 1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); 2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); 3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [20]). Table 4 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE). Our proposed formula $f$ consistently achieves the lowest EE and FE. $f_{1}$ struggles with finite-answer tasks where small models have non-zero scores. $f_{2}$ inaccurately assumes perfect scores are attainable, ignoring data limitations and task ambiguities. The direct power law ($f_{3}$) fails to model the 0-1 metric range and the varying difficulty of improvement near random guess and saturation. TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping (difficulty distributions visualized in Appendix F).
5.3.3 Different Anchor Points in Interpolation Mapping
We test three configurations:
Table 5: Influence of anchor points in the mapping stage on prediction error (pp). Lower is better.
Method Overall Individual Task Sets
[…]
w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
w. OOD anchor 1.59 2.56 2.22 0.94 1.84 1.04 2.56 1.86 1.18 1.04
w. ID anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
Table 5 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy. This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models. Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models. Additional ablations on the interpolation method (Sec. E.2) confirms quartic functions are suitable.
@@ 6 Conclusion and Discussion
In this work, we introduce a novel downstream performance scaling framework including (1) the COD method that effectively models the underlying diverse scaling patterns of evaluation tasks; (2) a scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions. We discuss the limitations of the COD method in Appendix G. Looking forward, our approach can be further expanded across model architectures, training methods, and evaluation set types, while extending this framework to address chain-of-thought reasoning patterns offer promising avenues for future research.
Boarder Impact
The LLMs performance scaling prediction method proposed in this work can be used to guide decision-making for training LLMs and to monitor whether the LLM training process aligns with expectations. Accurate prediction allows for better planning of compute and human resources, thereby playing a valuable role in scientifically advancing the training of larger LLMs. However, if the prediction results are inaccurate, it may lead to outcomes that do not meet expectations for the invested resources, potentially resulting in resource waste or economic losses.
@@ A.1 Improved MeanShift Algorithm
We iteratively apply the MeanShift algorithm with a predefined cluster radius $R$ and a minimum cluster size $K$. In each iteration, for the clustered samples, we examine whether the distance between each sample and its cluster center exceeds R, and relabel those samples that exceed this threshold as unclustered. For clusters containing fewer than K samples, we mark all samples in these clusters as unclustered. At the end of each iteration, we incorporate both the outliers from MeanShift and our marked unclustered samples into the next round of clustering, continuing this process until no further changes occur in sample labels. We present the pseudocode in algorithm 1.
@@
1:Initialize all labels in the evaluation set to $-1$
@@
3:     Perform MeanShift clustering with radius $R$ on all samples labeled $-1$
@@
12:         if number of samples in cluster $<K$ then
13:              Reset all samples in this cluster to $-1$
@@ Appendix B Proof of Theorem 1
We use lemma 1 to derive the scaling law for downstream task performance (theorem A1).
@@
$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\mu_{a}).$ (4)
@@
$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}\right)+o(\mu_{a})$
$=$ $\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o(\mu_{a})$
$=$ $\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\mu_{a}).$
@@
$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\mu_{a}).$ (7)
@@
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
  1. 1.
    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
  2. 2.
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
  3. 3.
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (9)
where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
@@
We first use Assumption 3 to establish the relationship between model passrate and loss on a task:
$-\log(p_{\mathrm{ans}})$ $=-\log\left(\prod_{t\in\mathrm{ans}}p(t)\right)=-\sum_{t\in\mathrm{ans}}\log(p_{t})=\mathrm{loss}_{\mathrm{ans}}.$ (10)
@@
$\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ $=\mathbb{E}_{p}[\exp(-\mathrm{loss}_{\mathrm{ans}}(C))]=\frac{1}{n}\sum_{(q,\mathrm{ans})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C)).$ (11)
@@
We can adopt lemma 1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
@@
$\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\exp(-\mathrm{loss}_{\mathrm{ans}}(C))=$ $\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,\mathrm{ans}_{t})\in P}\mathrm{loss}_{\mathrm{ans}}(C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (12)
$=$ $\exp{(-aC^{-b}-c)}+\frac{\sigma^{2}}{2\mu}+o(\mu)$ (13)

### v3

@@ 5.2 Prediction Experiments
We compare COD against four representative prediction methods:
@@
    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws [43].
@@
    Passrate-compute prediction [1, 23]: A variant of end-to-end method, which estimates large model passrates from smaller model passrates. We conduct 100 trials per evaluation set for smaller models to enhance reliability and report absolute prediction error on the passrate metric.
  4. 4.
    BNSL [7]: Another variant of end-to-end method, which decomposes the end-to-end mapping into a multi-segment power-law framework.
Table 1: Absolute prediction error (%) on evaluation sets for predicting the performance of the 70B model. Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red). $\downarrow$ indicates lower is better.
@@
    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within all clusters, and aggregates metrics, bypassing the subset-to-full mapping stage.
@@
Comparative results are shown in Tab.˜1. Prediction accuracy is measured by the absolute error between predicted and actual performance. We report mean and max prediction errors across all evaluation sets, as well as errors for individual sets. Our complete COD approach significantly outperforms existing methods in both mean (1.68%) and maximum (2.38%) prediction errors, offering reliable guidance for large model training. While baselines perform well on some sets, their substantial errors on others undermine overall credibility.
Figure 4: Performance-compute relationship for different prediction methods on eight evaluation sets.
figure˜4 visualizes the performance-compute relationships. The COD method does not merely extend the existing scaling trend; instead, it effectively predicts whether growth slowdown will occur subsequently and enables better estimation of the magnitude of curve bending. On BBH, while all methods give similar final estimates, end-to-end and loss-intermediate methods fit small model data poorly. COD reveals a more complex, better-fitted multi-phase trajectory. On MATH and MMLU-pro, where predicting accelerated growth versus plateaus is crucial, the loss-intermediate method underestimates model ceilings, and the end-to-end method exceeds 3% error. COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets and capture diminishing returns in saturated sets.
To further validate the generalizability of the COD method, we examine its predictive capabilities across different model scales and for models after continual training. In addition to the 70B model, the COD method also demonstrates effectiveness in predicting the capabilities of the medium-size models, with experimental details and corresponding conclusions available in Appendix E. Beyond this, the method is further shown to be capable of predicting the performance of LLMs following a continual training phase on high-quality data. To validate this, we conducted dedicated experiments and in-depth analyses on continually trained LLMs, with results documented in Appendix F.
@@ 5.3 Ablation Study
In this section, we conduct ablation experiments on the clustering method and extrapolation function, while in the Appendix C we discuss ablation studies on the selection criteria for extrapolatable subsets, interpolation mapping methods, and the impact of anchor point selection on predictions.
@@ 5.3.1 Comparison of Clustering Methods
Dataset K-Means DBScan MeanShift Improved-KMeans Improved-MeanShift
IAD OR $E$ FE IAD OR $E$ FE IAD OR $E$ FE IAD OR $E$ FE IAD OR $E$ FE
MMLU-pro 0.32 - 3.69 3.69 0.42 0.56 3.72 3.69 0.29 0.39 3.15 3.08 0.16 2.85 0.56 0.61 0.22 4.40 1.27 1.35
GSM8k 0.22 - 0.01 0.00 0.51 0.53 4.08 4.12 0.29 0.61 0.67 0.74 0.13 2.73 3.92 4.08 0.19 4.93 2.20 2.23
MATH 0.22 - 2.62 2.34 0.48 0.68 4.38 4.16 0.21 1.44 2.55 2.26 0.09 2.22 0.81 0.51 0.15 2.66 1.14 1.28
BBH 0.63 - 8.16 8.99 0.71 18.92 3.53 4.36 0.27 20.72 2.12 0.65 0.20 37.23 0.02 2.17 0.21 33.58 0.29 1.77
TriviaQA 0.44 - 2.97 2.46 0.70 6.38 1.11 0.81 0.25 6.77 3.64 4.90 0.12 11.97 1.18 1.12 0.19 11.54 1.58 1.64
AGIEval 0.46 - 2.61 2.68 0.56 3.67 6.43 6.27 0.29 2.99 2.63 3.23 0.15 7.60 5.96 5.56 0.21 11.50 1.11 2.38
DROP 0.56 - 1.66 1.64 0.67 11.08 3.03 2.66 0.25 11.81 4.18 4.00 0.14 21.42 3.99 5.24 0.20 19.88 0.26 0.23
MBPP 0.34 - 2.53 2.67 0.51 12.80 1.57 1.41 0.22 15.60 2.40 1.22 0.17 19.40 2.39 3.25 0.17 21.60 1.56 2.19
Mean - - 3.62 3.76 - - 3.93 4.08 - - 2.12 1.68 - - 1.33 1.84 - - 1.23 1.66
Max - - 8.16 8.99 - - 4.38 4.36 - - 3.15 3.08 - - 3.92 4.08 - - 2.20 2.23
Tab.˜2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints. The results also confirm these methods lead to lower EE and FE. Although Improved-KMeans has the best IAD, its GSM8k prediction is poor. This is likely because K-Means requires pre-specifying $k$, and our search for $k$ can be unstable, leading to large errors on some sets. In contrast, our Improved-MeanShift, which automatically determines $k$ based on distance constraints, offers more stable clustering and the lowest maximum prediction error.
@@ 5.3.2 Different Extrapolation Formulas
Table 3: Ablation study on extrapolation formulas. $E$, TR, FE shown for BBH, MATH, MMLU-pro.
@@
We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: (1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); (2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); (3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [23]). Tab.˜3 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE). Our proposed formula $f$ consistently achieves the lowest EE and FE. $f_{1}$ struggles with finite-answer tasks where small models have non-zero scores. $f_{2}$ inaccurately assumes perfect scores are attainable, ignoring data limitations and task ambiguities. The direct power law ($f_{3}$) fails to model the 0-1 metric range and the varying difficulty of improvement near random guess and saturation. TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping.
@@ 6 Conclusion and Discussion
In this work, we introduce a novel downstream performance scaling framework including (1) the COD method that effectively models the underlying diverse scaling patterns of evaluation tasks; (2) a scaling law for downstream task performance that provides a fitting formula for performance-compute extrapolation; and (3) a systematic methodology for identifying and leveraging predictable subset that provides a robust intermediate metric for accurate full-set performance predictions. We discuss the limitations and future works in Appendix I.
Ethics Statement
We have read and adhered to the ICLR Code of Ethics. This work proposes a computational framework to enable more efficient resource allocation in the training of LLMs. The research is methodological in nature and aims to support more sustainable and responsible practices within the field of AI development.
We provide a detailed account of our methods, theoretical proofs, and experimental settings. We openly discuss the limitations of our framework in Appendix I. This study does not involve human subjects or the use of sensitive personal data.
We utilized language models at the writing level, including checking for grammatical errors in the article and modifying expressions. The use of language models had no impact on the article’s innovative contributions, experiments, or analytical perspectives.
Reproducibility Statement
We have made extensive efforts to ensure the reproducibility of our work. The core methodology, the Clustering-On-Difficulty (COD) framework, is detailed in Sec.˜4. The improved MeanShift clustering algorithm is described in Sec.˜4.1 , with full pseudocode provided in Appendix A.1 (Algorithm 1). Our performance scaling law (Theorem 1) is presented in Sec.˜4.2 , with a complete proof available in appendix˜B. We discuss the additional computational cost of the COD method in Appendix H.
Our experimental setup, including model architectures, training data philosophy, and hyperparameters, is thoroughly documented in Sec.˜5.1 and Appendix D , with specific model configurations listed in Tab.˜A4. The evaluation benchmarks, protocols, and few-shot settings are described in Sec.˜5.1 and summarized in Tab.˜A5. Extensive ablation studies validating our component choices—including extrapolation formulas (Sec.˜5.3.2), clustering algorithms (Appendix 5.3.1), interpolation methods (Appendix C.2), and criteria for filtering clusters (Appendix C.1)—are provided to support our findings. We visualize the task difficulty distribution for each evalset in Appendix G.
@@ A.1 Improved MeanShift Algorithm
We iteratively apply the MeanShift algorithm with a predefined cluster radius $R$ and a minimum cluster size $K$. In each iteration, for the clustered samples, we examine whether the distance between each sample and its cluster center exceeds R, and relabel those samples that exceed this threshold as unclustered. For clusters containing fewer than K samples, we mark all samples in these clusters as unclustered. At the end of each iteration, we incorporate both the outliers from MeanShift and our marked unclustered samples into the next round of clustering, continuing this process until no further changes occur in sample labels. We present the pseudocode in algorithm˜1.
@@
1:Initialize all labels in the evaluation set to 1-1
@@
3:  Perform MeanShift clustering with radius $R$ on all samples labeled 1-1
@@
12:   if number of samples in cluster <K<K then
13:     Reset all samples in this cluster to 1-1
@@
In the experiment, $K$ is empirically set to $10$, which has been verified through extensive experiments to be a reasonable and robust value. The fixed radius $R$ depends only on the value range and dimensionality (i.e., the number of predictive models) of the difficulty feature vectors. In our experiments, we fixed $R=0.3$ for all evaluation sets to ensure consistent overall clustering performance. For experiments with different numbers of small models used for fitting, the value of $R$ needs to be adaptively adjusted due to changes in the dimensionality of the feature vector. For instance, in the experiment of predicting the 12B-parameter model, $R$ was empirically reduced from $0.3$ to $0.2$ to ensure that the intra-class error is reasonable.
@@ Appendix B Proof of Theorem 1
We use lemma˜1 to derive the scaling law for downstream task performance (theorem˜1).
@@
$\Delta$ $=\mu_{a}-\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}x_{i}-\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=\frac{\sigma^{2}}{2\mu_{a}}+o(\sigma^{2}).$ (6)
@@
$\frac{1}{n}\sum_{i=1}^{n}\log x_{i}=$ $\log\mu+\frac{1}{n}\sum_{i=1}^{n}\left(\frac{(x_{i}-\mu_{a})}{\mu_{a}}-\frac{(x_{i}-\mu_{a})^{2}}{2\mu_{a}^{2}}+o\left((x_{i}-\mu_{a})^{2}\right)\right)$
$=$ $\log\mu+\frac{1}{\mu}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}-\mu_{a}\right)}_{\text{equal to 0}}+\frac{1}{2\mu_{a}^{2}}\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)}_{\sigma^{2}}+o\left(\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\mu_{a})^{2}\right)$
$=$ $\log\mu-\frac{\sigma^{2}}{2\mu^{2}}+o(\sigma^{2}).$
@@
$\mu_{a}-\mu_{g}=\mu_{a}\left(1-\exp{\left(-\frac{\sigma^{2}}{2\mu^{2}}\right)}\right)+o(\sigma^{2}).$ (9)
@@
Consider a language model $M_{C}$ trained with compute budget $C$ and a set of downstream tasks $\mathcal{P}$. Under the following assumptions: Assumption 1 (Power-law scaling of answer loss): the expected answer loss follows:
$L_{\mathcal{P}}(C):=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]=\alpha C^{-\beta}+\gamma$ (11)
where $\alpha,\beta,\gamma>0$ are task-specific constants, with $\gamma$ representing the irreducible loss.
Assumption 2 (Unique deterministic answers): Each question has a unique deterministic answer. The model receives score $1$ if and only if $M_{C}$ outputs $a_{\text{true}}$, and 0 otherwise.
Assumption 3 (Accuracy decomposition): The expected accuracy decomposes as:
$\mathbb{E}_{T\sim\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\cdot\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[p(a_{\text{true}}|q,M_{C})],$ (12)
where $g\in[0,1]$ is the random guessing baseline.
Then, the expected accuracy on task set $\mathcal{P}$ can be modeled as:
$\mathbb{E}_{\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\left(\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}\right)+o\left(\sigma_{L}^{2}(C)\right)$ (13)
where $\mu_{L}(C)=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the mean loss and $\sigma_{L}^{2}(C)=\mathrm{Var}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the loss variance across the task set.
@@
For a task $T=(q,a_{\text{true}})\in\mathcal{P}$, under assumption 2, $a_{\text{true}}$ is deterministic and unique, thus $p(a_{\text{true}}|q,M_{C})$ can be decomposed into token-wise auto-regressive loss.
$-\log(p(a_{\text{true}}|q,M_{C}))$ $=-\log\left(\prod_{i=1}^{k}p(t_{i}|q,t_{<i};M_{C})\right)$ (14)
$=-\sum_{i=1}^{k}\log\left(p(t_{i}|q,t_{<i};M_{C})\right)$ (15)
$=L(q,a_{\text{true}};C).$ (16)
@@
$\mathbb{E}_{p}[p(a_{\text{true}}|q,M_{C})]$ $=\mathbb{E}_{p}[\exp(-L(q,a_{\text{true}};C)]=\frac{1}{n}\sum_{(q,a_{\text{true}})\in P}\exp(-L(q,a_{\text{true}};C)).$ (17)
@@
We can adopt lemma˜1 to switch from arithmetic mean to geometric mean of $\mathrm{loss}$, and apply the power law assumption 1.
@@
$\frac{1}{n}\sum_{(q,a_{\text{true}})\in P}\exp(-L(q,a_{\text{true}};C))=$ $\exp{\underbrace{\left(-\frac{1}{n}\sum_{(q,a_{\text{true}})\in P}L(q,a_{\text{true}};C)\right)}_{\text{use loss scaling law}}}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}+o\left(\sigma_{L}^{2}(C)\right)$ (18)
$=$ $\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}+o\left(\sigma_{L}^{2}(C)\right)$ (19)